# Constraint-Aware Reinforcement Learning via Adaptive Action Scaling

### v1

@@ V Experiments
Environments: We evaluate our method on locomotion tasks from the Safety Gym benchmark [21], namely Ant, Walker2d, Swimmer, HalfCheetah, and Humanoid. In these velocity tasks, a safety cost is incurred whenever the center-of-mass speed exceeds a predefined threshold. Because the cost signal is sparse and triggered only by such threshold violations, these environments provide a challenging setting for safe RL, while naturally aligning with our regulator’s goal of attenuating large actions that are most likely to induce violations. We further evaluate on BiGlucose and the Continuous Stirred Tank Reactor (CSTR) from the Safety-Critical Systems  [22].
Baselines: We compare our regulator against five state-of-the-art Safe RL baselines. PPO-PID [17] augments PPO with a PID-controlled Lagrangian multiplier to mitigate instabilities commonly observed in dual updates during constrained optimization. Simmer [19] augments PPO with a safety state that tracks the remaining safety budget. Safety Editor [24] uses two SAC agents, one for maximizing the reward and another for editing unsafe actions. RESPO [25] estimates reachability sets and constrains policy to remain within safe regions. SRCPO [26] formulates a bi-level constrained optimization using spectral risk measures to achieve a near-zero constraint violation rate while maximizing reward.
Metrics: Similar to prior safe RL studies, we report returns and cumulative costs as in [23, 25, 32], and additionally follow [23] in using the return-to-cost (RC) ratio to capture the trade-off between task performance and safety. Specifically, we measure (i) episodic return, (ii) cumulative cost during training, which reflects the total number of constraint violations and, in sparse-cost settings such as Safety Gym velocity tasks, implicitly captures violation frequency, and (iii) the RC ratio, defined as the total return divided by cumulative cost. For visualization, we plot the return divided by the logarithm of the accumulative cost, which improves interpretability of the safety-performance trade-off.
V-A Comparison Against Baselines:
Safety Gym Results
@@ V-A Comparison Against Baselines:
PPO-PID [17] 27.33 3.04 0.32 3.76 7.13
SIMMER [19] 26.14 6.09 1.21 3.59 7.60
SEditor [24] 3.34 5.64 -0.10 1.95 18.52
RESPO [25] 23.97 2.89 0.17 204.40 12.73
SRCPO [26] -0.11 0.02 2.90 0.86 0.90
@@
PPO-PID [17] 39.29 28.39 21.31 19.11 5.46
SIMMER [19] 41.88 13.70 54.22 14.14 3.64
SEditor [24] 16.19 1.67 1.77 0.48 1.88
RESPO [25] 8.36 1.65 27.39 2.03 1.29
SRCPO [26] 23.24 23.55 37.10 5.07 2.82
TABLE I: Relative return improvement (↑) and relative cumulative cost (↓) of SAC-Reg compared to baselines across locomotion tasks. SAC-Reg consistently achieves higher returns and lower cumulative costs than prior safe RL methods across all environments.
@@
PPO-PID [17] 18.49 3.05 0.17 3.87 4.51
SIMMER [19] 17.70 6.11 1.18 3.70 4.83
SEditor [24] 2.55 5.65 -0.20 2.02 12.22
RESPO [25] 16.26 2.90 3.59 209.23 8.30
SRCPO [26] -0.41 0.02 2.46 0.91 0.29
@@
PPO-PID [17] 60.14 19.77 34.06 126.18 5.20
SIMMER [19] 64.11 9.54 86.65 93.37 3.47
SEditor [24] 24.78 1.16 2.84 3.15 1.74
RESPO [25] 12.80 1.15 43.78 13.42 1.23
SRCPO [26] 35.57 16.39 59.28 33.50 2.69
TABLE II: Relative return improvement (↑) and relative cumulative cost (↓) of TD3-Reg compared to baselines across locomotion tasks.TD3-Reg demonstrates similar trends, outperforming baselines in return while maintaining substantially lower cumulative costs.
Tables I and II summarize results for both regulators against established baselines. Two metrics are reported: the relative return improvement,
@@
Experiments on Safety Critical Systems
Figure 4: Performance comparison on the BiGlucose and CSTR environments. Our method (SAC-Reg) achieves high return with low cost, yielding the best return-to-cost ratio throughout training.
@@
We evaluate our approach on two continuous-control environments: BiGlucose and the Continuous Stirred Tank Reactor (CSTR) from [22], which capture biomedical and chemical process dynamics. BiGlucose models blood glucose regulation with insulin and glucagon injections under delayed, partially observable dynamics, requiring glucose levels to stay within physiological bounds. CSTR simulates nonlinear reactor dynamics where unsafe control can cause hazardous runaway reactions. In both cases, we modify the environments to provide a continuous cost signal by logging violation magnitudes, instead of terminating episodes on constraint breaches. For full specifications, see [22].
Figure 4 compares our SAC-Reg method against baselines (SEditor, SIMMER, SRCPO, and PPO-PID). All methods are trained for the same number of environment steps as in [22]; RESPO is excluded since it takes significantly more steps to converge. Across both tasks, SAC-Reg consistently achieves higher returns with fewer cumulative constraint violations, yielding superior safety–performance trade-offs. In BiGlucose, baselines such as SIMMER and SRCPO quickly accumulate costs despite improving return, while SEditor remains return-limited. In CSTR, only our method manages both safety and performance, as others either accumulate high violations or fail to learn.
Ablation Study on $\lambda$ and $\beta$
We conduct an ablation study in the Ant environment to evaluate the sensitivity of our regulator framework to the hyperparameters $\lambda$ and $\beta$, which control the trade-off between action retention and cost suppression. When varying $\lambda$ over the range $\{1\times 10^{-5},\ 0.0015,\ 0.05,\ 0.25,\ 1.0\}$, we find that smaller values lead to significantly lower cumulative costs. In particular, $\lambda=0.0015$ achieves the best balance between constraint satisfaction and task performance. Higher values of $\lambda$ result in larger action magnitudes and consequently higher constraint violations. Similarly, varying $\beta$ over $\{5,\ 10,\ 15,\ 30,\ 50\}$ shows that $\beta=10$ achieves the best overall safety-performance balance, minimizing constraint violations while maintaining high return. These results, as shown in Fig. 5, highlight the importance of properly tuning the regulator’s loss coefficients to achieve optimal return-to-cost behavior.
Element-wise vs. Scalar Regulation
To evaluate the impact of element-wise action regulation, we conducted an ablation study comparing our full regulator with a simplified variant that uses a single scalar value to uniformly scale all action dimensions. Figure 6 presents results across all Safety Gymnasium locomotion tasks. While the scalar variant achieves comparable performance in most environments, it fails to converge in the high-dimensional Humanoid task. This suggests that element-wise scaling is particularly important in complex control settings, where individual action dimensions exhibit distinct risk profiles. Fine-grained modulation allows the regulator to target risky joints more precisely, improving both safety and learning stability.
@@
TABLE III: Training performance under different levels of injected Gaussian noise ($\sigma=0.00,0.025,0.05,0.10$) in observations and actions. Values show episode return and cumulative cost at checkpoints. Our regulator maintains bounded costs across noise levels, demonstrating robustness relevant for sim-to-real transfer.
Robustness and Sim-to-Real Transfer
To approximate uncertainties encountered on physical robots, we inject Gaussian noise into both observations and actions during training, modeling sensor measurement errors and actuator execution noise. Agents are trained with noise levels ($\sigma=0,0.025,0.05,0.10$), and the resulting training performance is summarized in Table III. Across all noise settings, our regulator achieves strong returns while keeping cumulative costs bounded. Even under the highest noise level ($\sigma=0.10$), performance remains stable, highlighting robustness to sensing and actuation imperfections and supporting the method’s potential for sim-to-real transfer.
VI Conclusion

### v2

@@ 5 Experiments
Environments: We evaluate our method on locomotion tasks from the Safety Gym benchmark [ji2023safety], namely Ant, Walker2d, Swimmer, HalfCheetah, and Humanoid. In these velocity tasks, a safety cost is incurred whenever the center-of-mass speed exceeds a predefined threshold. Because the cost signal is sparse and triggered only by such threshold violations, these environments provide a challenging setting for safe RL, while naturally aligning with our regulator’s goal of attenuating large actions that are most likely to induce violations.
Baselines: We compare our regulator against five state-of-the-art Safe RL baselines. PPO-PID [stooke2020responsive] augments PPO with a PID-controlled Lagrangian multiplier to mitigate instabilities commonly observed in dual updates during constrained optimization. Simmer [sootla2022enhancing] augments PPO with a safety state that tracks the remaining safety budget. Safety Editor [yu2022towards] uses two SAC agents, one for maximizing the reward and another for editing unsafe actions. RESPO [ganai2023iterative] estimates reachability sets and constrains policy to remain within safe regions. SRCPO [kim2024spectral] formulates a bi-level constrained optimization using spectral risk measures to achieve a near-zero constraint violation rate while maximizing reward.
Metrics: Similar to prior safe RL studies, we report returns and cumulative costs as in [thananjeyan2021recovery, ganai2023iterative, goodall2024leveraging], and follow [thananjeyan2021recovery] in using the return-to-cost (RC) ratio to capture the trade-off between task performance and safety. We measure (i) episodic return, (ii) cumulative cost during training, which reflects the total number of constraint violations and, in sparse-cost settings such as Safety Gym velocity tasks, implicitly captures violation frequency, and (iii) the RC ratio, defined as the total return divided by cumulative cost. For visualization, we plot the return divided by the logarithm of the accumulative cost, which improves interpretability of the safety-performance trade-off.
5.1 Comparison Against Baselines:
Safety Gym Results.
@@ 5.1 Comparison Against Baselines:
PPO-PID [stooke2020responsive] 27.33 3.04 0.32 3.76 7.13
SIMMER [sootla2022enhancing] 26.14 6.09 1.21 3.59 7.60
SEditor [yu2022towards] 3.34 5.64 -0.10 1.95 18.52
RESPO [ganai2023iterative] 23.97 2.89 0.17 204.40 12.73
SRCPO [kim2024spectral] -0.11 0.02 2.90 0.86 0.90
@@
PPO-PID [stooke2020responsive] 39.29 28.39 21.31 19.11 5.46
SIMMER [sootla2022enhancing] 41.88 13.70 54.22 14.14 3.64
SEditor [yu2022towards] 16.19 1.67 1.77 0.48 1.88
RESPO [ganai2023iterative] 8.36 1.65 27.39 2.03 1.29
SRCPO [kim2024spectral] 23.24 23.55 37.10 5.07 2.82
Table 1: Relative return improvement (↑) and relative cumulative cost (↓) of SAC-Reg compared to baselines across locomotion tasks. SAC-Reg consistently achieves higher returns and lower cumulative costs than prior safe RL methods across all environments.
@@
PPO-PID [stooke2020responsive] 18.49 3.05 0.17 3.87 4.51
SIMMER [sootla2022enhancing] 17.70 6.11 1.18 3.70 4.83
SEditor [yu2022towards] 2.55 5.65 -0.20 2.02 12.22
RESPO [ganai2023iterative] 16.26 2.90 3.59 209.23 8.30
SRCPO [kim2024spectral] -0.41 0.02 2.46 0.91 0.29
@@
PPO-PID [stooke2020responsive] 60.14 19.77 34.06 126.18 5.20
SIMMER [sootla2022enhancing] 64.11 9.54 86.65 93.37 3.47
SEditor [yu2022towards] 24.78 1.16 2.84 3.15 1.74
RESPO [ganai2023iterative] 12.80 1.15 43.78 13.42 1.23
SRCPO [kim2024spectral] 35.57 16.39 59.28 33.50 2.69
Table 2: Relative return improvement (↑) and relative cumulative cost (↓) of TD3-Reg compared to baselines across locomotion tasks.TD3-Reg demonstrates similar trends, outperforming baselines in return while maintaining substantially lower cumulative costs.
Tables 1 and 2 summarize results for both regulators against established baselines. Two metrics are reported: the relative return improvement,
@@
Ablation Study on $\lambda$ and $\beta$. We conduct an ablation study in the Ant environment to evaluate the sensitivity of our regulator framework to the hyperparameters $\lambda$ and $\beta$, which control the trade-off between action retention and cost suppression. When varying $\lambda$ over the range $\{1\times 10^{-5},\ 0.0015,\ 0.05,\ 0.25,\ 1.0\}$, we find that smaller values lead to significantly lower cumulative costs. In particular, $\lambda=0.0015$ achieves the best balance between constraint satisfaction and task performance. Higher values of $\lambda$ result in larger action magnitudes and consequently higher constraint violations. Similarly, varying $\beta$ over $\{5,\ 10,\ 15,\ 30,\ 50\}$ shows that $\beta=10$ achieves the best overall safety-performance balance, minimizing constraint violations while maintaining high return. These results, as shown in Fig. 4, highlight the importance of properly tuning the regulator’s loss coefficients to achieve optimal return-to-cost behavior.
@@
Element-wise vs. Scalar Regulation: To evaluate the impact of element-wise action regulation, we conducted an ablation study comparing our full regulator with a simplified variant that uses a single scalar value to uniformly scale all action dimensions. Figure 5 presents results across all Safety Gymnasium locomotion tasks. While the scalar variant achieves comparable performance in most environments, it fails to converge in the high-dimensional Humanoid task. This suggests that element-wise scaling is particularly important in complex control settings, where individual action dimensions exhibit distinct risk profiles. Fine-grained modulation allows the regulator to target risky joints more precisely, improving both safety and learning stability.
@@
Table 3: Training performance under different levels of injected Gaussian noise ($\sigma=0.00,0.025,0.05,0.10$) in observations and actions. Values show episode return and cumulative cost at checkpoints. Our regulator maintains bounded costs across noise levels, demonstrating robustness relevant for sim-to-real transfer.
Robustness and Sim-to-Real Transfer. To approximate uncertainties encountered on physical robots, we inject Gaussian noise into both observations and actions during training, modeling sensor measurement errors and actuator execution noise. Agents are trained with noise levels ($\sigma=0,0.025,0.05,0.10$), and the resulting training performance is summarized in Table 3. Across all noise settings, our regulator achieves strong returns while keeping cumulative costs bounded. Even under the highest noise level ($\sigma=0.10$), performance remains stable, highlighting robustness to sensing and actuation imperfections and supporting the method’s potential for sim-to-real transfer.
6 Conclusion
@@
Acknowledgments
M. Dawood, U. Ahmed Siddiquie, S. Khorshidi, and M. Bennwitz are with the Humanoid Robots Lab and the Center for Robotics, University of Bonn, Germany. Additionally, M. Dawood and M. Bennwitz are with the Lamarr Institute for Machine Learning and Artificial Intelligence. This work has partially been funded by the German Federal Ministry of Research, Technology and Space (BMFTR) under the Robotics Institute Germany (RIG).